# Colimits of Heyting Algebras through Esakia Duality

### v1

@@ Abstract
In this note we generalize the construction, due to Ghilardi, of the free Heyting algebra generated by a finite distributive lattice, to the case of arbitrary distributive lattices. Categorically, this provides an explicit construction of a left adjoint to the inclusion of Heyting algebras in the category of distributive lattices This is shown to have several applications, both old and new, in the study of Heyting algebras: (1) it allows a more concrete description of colimits of Heyting algebras, as well as, via duality theory, limits of Esakia spaces, by knowing their description over distributive lattices and Priestley spaces; (2) It allows a direct proof of the amalgamation property for Heyting algebras, and of related facts; (3) it allows a proof of the fact that the category of Heyting algebras is co-distributive. We also study some generalizations and variations of this construction to different settings. First, we analyse some subvarieties of Heyting algebras – such as Boolean algebras, $\mathsf{KC}$ and $\mathsf{LC}$ algebras, and show how the construction can be adapted to this setting. Second, we study the relationship between the category of image-finite posets with p-morphisms and the category of posets with monotone maps, showing that a variation of the above ideas provides us with an appropriate general idea.
@@ 1 Introduction
One of the many attractive features of duality theory as it pertains to the theory of Boolean algebras and distributive lattices is that many constructions which are quite generic in algebraic terms (e.g., word constructions), become more concrete when put in dual terms. A prime example of this lies in the coproduct of Boolean algebras: in contrast to the abstract description of coproducts of algebras, given Boolean algebras $\mathcal{B}_{0}$ and $\mathcal{B}_{1}$, their coproduct $\mathcal{B}_{0}\oplus\mathcal{B}_{1}$ can be described as
$\mathsf{Clop}(X_{0}\times X_{1}),$
where $\mathcal{B}_{0}^{*}=X_{0}$ and $\mathcal{B}_{1}^{*}=X_{1}$ are the Stone duals of these algebras. This makes it possible to prove several general facts by simple observations – for instance, the fact that the canonical coprojections into a coproduct are injective follows from the fact that the projections of Stone spaces onto its factors are surjective. Similar facts hold regarding distributive lattices and their dual Priestley spaces.
At the root of this simplicity in describing coproducts, lies the fact that free algebras in these varieties admit rather explicit constructions. Indeed, the structure of free Boolean algebras and free distributive lattices on any number of generators is well-understood, and has been extensively exploited both for algebraic and set-theoretic purposes (see [21, Chapter 9-10] for an extensive discussion).
In sharp contrast to this, the structure of free Heyting algebras is much more difficult: already the free algebra on one generator is infinite, and the free algebra on two generators is notoriously difficult. As such, there has been considerable attention devoted to such objects. Notably, the construction of the $n$-universal model (due independently to Bellissima [3], Grigolia [18]) and Shehtman [27], allowed a more concrete description of the completely join-irreducible part of the free $n$-generated Heyting algebras; whilst the construction of the free $n$-generated Heyting algebra by Ghilardi [16] (which generalized previous work by Urquhart [29], and was in turn generalized for all finitely presented Heyting algebras by Butz [9]) and the associated ideas allowed the development of semantic proofs of several key facts regarding this category of algebras: the fact that free Heyting algebras are bi-Heyting algebras, the fact that the first-order theory of Heyting algebras has a model completion (due in its syntactic form to [24], and proven as such by Ghilardi and Zawadowski [17]), the fixed-point elimination theorem (proved in its syntactic form by Ruitenburg [23] and in its semantic form by Ghilardi and Santocanale [26]), amongst others. All of these make explicit and non-trivial use of the fact that the algebras of interest are finitely presented, and exploit the finitarity to work with discrete dualities at key steps. As such, a description of general free constructions of Heyting algebra seems to not have been addressed in the current literature.
From our point of view, the main hurdle in this is realizing the crucial point of finiteness: the above constructions by Ghilardi rely heavily on the properties of Birkhoff duality for finite distributive lattices, and it is not obvious how to generalize such facts to the infinite case. The purpose of this paper is to provide such a generalization. In place of Birkhoff duality we use Priestley duality, and show that this allows us to construct all free Heyting algebras in a way which generalizes existing work. This amounts essentially to describing a left adjoint to the inclusion of $\mathbf{HA}$, the category of Heyting algebras, in $\mathbf{DL}$, the category of distributive lattices.
Having such a description at hand, we then proceed to employ it to provide some results on the structure of Heyting algebras, some known and some new:
@@
    A description of free Heyting algebras on any number of generators is given;
@@
    A description of coproducts of Heyting algebras is given, and it is shown that the category of Heyting algebras is co-distributive;
@@
    A description of pushouts of Heyting algebras is given, and it is shown directly that the coprojections of Heyting algebras to the pushout are injective (yielding, as a corollary, the amalgamation property).
We then proceed to extend these results this in several directions, remaining in the realm of Heyting algebras and associated categories. We first show that this construction can be adapted when dealing with some subvarieties of Heyting algebras. Rather than providing a general theory, we focus on two important and illustrative cases: $\mathsf{KC}$ algebras and $\mathsf{LC}$ algebras. Our analysis of the latter also reveals a surprising property of Gödel logic which we were not aware before: that every formula in Gödel logic is equivalent to one of modal depth $2$. Finally, we look at the relationship between the category of posets with p-morphisms and the category of posets with monotone maps, and provide a left adjoint to the inclusion which is heavily inspired by the above constructions. We briefly discuss some coalgebraic consequences of this, which we plan to elaborate on in joint work with Nick Bezhanishvili [2].
The structure of the note is as follows: in Section 2 we recall some basic preliminaries, and fix some notation moving forward. Section 3 contains the main technical tools, in the form of the description of the free Heyting algebra generated by a distributive lattice preserving a given set of relative pseudocomplements; in Section 4 we showcase some basic applications of the aforementioned theory to the study of Heyting algebras: some well-known results, like the amalgamation theorem, are shown here through direct methods, as well as some seemingly new results, like the co-distributivity of the category of Heyting algebras. In Section 5 we look at the case of subvarieties of Heyting algebras and illustrate how these constructions can be fruitfully analysed there. In Section 6 we look at the corresponding relationship between the category of posets with p-morphisms and the category of posets with monotone maps. We conclude in Section 7 with some outlines of future work.
@@ 2 Preliminaries
2.1 Algebraic Background
We assume the reader is familiar with the basic theory of Boolean algebras and Distributive lattices (see for example [12]). Throughout we will use the term “distributive lattices” for bounded distributive lattices, i.e., our lattices will always have a bottom element $(0)$ and a top element $(1)$. We also assume the reader is familiar with Heyting algebras and their elementary theory (see e.g. [15]).
We recall the following constructions, which are part of the folklore of the subject (see [28] for an in-depth discussion):
@@ 2.1 Algebraic Background
Let $X$ be an arbitrary set. Let $\mathbf{B}(X)$ be the Boolean algebra dual to the Stone space
$F(X)\coloneqq\mathbf{2}^{X}$
where $\mathbf{2}$ is the two element Stone space, and $F(X)$ is given the product topology.
Similarly, let $\mathbf{D}(X)$ be the distributive lattice dual to the Priestley space
$F_{D}(X)\coloneqq\mathbf{2}^{X}$
where $\mathbf{2}$ is this time the poset $0<1$, and $F_{D}(X)$ is given the product topology and the product ordering.
We then have the following well-known fact:
Proposition 2.
Let $X$ be an arbitrary set. Then:
  1. 1.
    $\mathbf{B}(X)$ is the free Boolean algebra on $X$ many generators;
  2. 2.
    $\mathbf{D}(X)$ is the free distributive lattice on $X$ many generators.
Given $\mathcal{D}$ a distributive lattice, and $A\subseteq D$, we write $\mathsf{Fil}(A)$ for the filter generatedd by $A$, and similarly, $\mathsf{Id}(A)$ for the ideal generated by $A$. We also recall that the following prime filter theorem holds for distributive lattices:
@@
2.2 Duality Theory
Throughout we assume the reader is familiar with Stone duality, as covered for example in [12]. Given a distributive lattice, we denote the representation map as
$\varphi:\mathcal{D}\to\mathsf{ClopUp}(X_{D})$
where $\varphi(a)=\{x\in X_{D}:a\in x\}$.
We also note that for our purposes a given ordered topological space $(X,\leq)$ is called a Priestley space if:
  • $(X,\tau)$ is a compact topological space.
  • For each $x,y\in X$, if $x\nleq y$ then there is a clopen upset $U$ such that $x\in U$ and $y\notin U$.
We say that a Priestley space $X$ is an Esakia space if whenever $U\subseteq X$ is a clopen set, then ${\downarrow}U$ is likewise clopen.
A fact that will often use below is that if $(X,\leq)$ is a Priestley space, and $Y\subseteq X$ is a closed subset, then $(Y,\leq_{\restriction Y})$ is a Priestley space as well. We also note the following important, though easy to prove, fact:
@@ 2.2 Duality Theory
Let $(\mathcal{D}_{n},\iota_{i,j})$ be a chain of distributive lattices, connected by embeddings $i_{i,j}:\mathcal{D}_{i}\to\mathcal{D}_{j}$ for $i<j$ satisfying the compatibility laws. Let $(X_{n},p_{i,j})$ be the inverse system of Priestley spaces one obtains by dualising all arrows. Then the directed union of the directed chain, $\mathcal{D}_{\omega}$, is dual to the projective limit $X_{\omega}$ of the inverse system. Moreover, if each $X_{n}$ is an Esakia space, and each $p_{i,j}$ is a p-morphism, then $X_{\omega}$ is likewise an Esakia space.
We also recall the well-known construction of the Vietoris hyperspace, due to Leopold Vietoris [30]: let $(X,\leq)$ be a Priestley space. Let $V(X)$ be the set of all closed subsets of $X$. We give this set a topology by considering a subbasis consisting, for $U,V\subseteq X$ clopen sets:
@@
This is sometimes called the “hit-and-miss topology”, and the resulting space is called the Vietoris hyperspace. On this space we define an order relation $C\preceq D$ if and ony if [^1] The reason this order is reverse inclusion has to with the choice to use upsets rather than downsets; in [16], the author uses downsets, which is why the swap no reversion is needed. $D\subseteq C$. Then we have the following:
@@
The fact that $V(X)$ is compact is a standard fact (see e.g. [14]), but we prove it here for completeness. assume that $V(X)=\bigcup_{i\in I}[U_{i}]\cup\bigcup_{j\in J}\langle V_{j}\rangle$ is a cover by clopen sets. Now consider $C=X-\bigcup_{i\in I}V_{i}$. If $C$ is empty, then the sets $V_{i}$ cover $X$ and so finitely many of them cover $X$, say $V_{i_{0}},...,V_{i_{n}}$. Then if $A\in V(X)$, then $A$ is non-empty and so contains some $x$ which must lie in some $V_{i_{k}}$ for $k\in\{1,...,n\}$, and so $A\in\langle V_{i_{k}}\rangle$. So the sets $\{\langle V_{i_{k}}\rangle:k\in\{1,...,n\}\}$ form a finite subcover.

### v3

# references.bib

@@ Abstract
In this paper we provide an explicit construction of the left adjoint to the inclusion of Heyting algebras in the category of distributive lattices using duality, generalizing a technique due to Ghilardi.This is employed to give a new concrete description of colimits and free Heyting algebras. Using such tools, we prove some results about the category of Heyting algebra. These are both known, concerning the amalgamation property and related facts; and new, such as the fact that the category of Heyting algebras is co-distributive. We also study some generalizations and variations of this construction to different related settings. We analyse some subvarieties of Heyting algebras – such as Boolean algebras, $\mathsf{KC}$ and $\mathsf{LC}$ algebras, and show how the construction can be adapted to such cases. Second, we study the relationship between the category of image-finite posets with p-morphisms and the category of posets with monotone maps, and using a variation of the above ideas, provide an adjunction between such categories.
@@ 1 Introduction
In algebraic settings, many problems boil down to the analysis of free objects: free groups, free modules over a ring, polynomial rings, amongst others. This ubiquity stems from the fact that understanding free objects of this kind allows a very fine combinatorial and categorical understanding of the structure of the classes of algebras of interest. As a result, they can encode a lot of the axiomatic structure of such algebras, as well as provide access to key theorems concerning finitely presented, finitely generated, or projective algebras, all of which intimately depend on free ones. Such algebras are always possible to describe using “word constructions”, or equivalently, by computing colimits of the algebras in terms of generators and relations, but such representations are often too abstract to yield useful results. As such, making them more concrete presents an attractive and often important problem.
In some algebraic settings, this concreteness can be obtained by duality theory. In its origins, in Stone’s celebrated theorem [Stone1936], this ‘duality’ associates to each given Boolean algebra $\mathcal{B}$ a unique compact, Hausdorff and zero-dimensional space $X_{B}$, corresponding to the prime spectrum of the Boolean algebra. The existence of such a duality makes several constructions much simpler: for instance, to calculate coproducts of Boolean algebras, one need only describe products of Stone spaces, which turn out to be exactly the usual topological products. Since arbitary free algebras can be computed as coproducts of $1$-generated algebras, this also allows an easy computation of free Boolean algebras.
A similar and more general situation holds concerning distributive lattices and their prime spectra. Here, with a given distributive lattice $\mathcal{D}$, one takes the prime spectrum $X_{D}$ equipped with an order $\leq_{D}$, such that the pair $(X_{D},\leq_{D})$ forms a so-called Priestley space [Priestley1970]. Here again the calculation of coproducts simplifies, dually, to the calculation of products of Priestley spaces, which are simple set-theoretic constructions.
The case of interest in this paper is that of Heyting algebras. These are objects in the signature $(H,\wedge,\vee,\rightarrow,0,1)$, where $(H,\wedge,\vee,0,1)$ is a bounded lattice, and where the meet operation $\wedge$ has an adjoint, i.e., for $a,b,c\in H$,
$a\wedge b\leq c\iff a\leq b\rightarrow c.$
These structures have been studied extensively in several different mathematical contexts, where they arise naturally: 1) they are algebraic models for intuitionistic logic [Chagrov1997-cr]; 2) all lattices of open sets of topological spaces are Heyting algebras [McKinsey1944]; 3) locales are complete Heyting algebras [PP2012]; 4) In categorical logic they arise as the decategorification of cartesian closed categories; 5) In topos theory they arise as the lattices of subobjects of toposes; amongst others. Following the work of Esakia [Esakiach2019HeyAlg], one can also represent the spectrum of a Heyting algebra in a way that accounts for the implication connective using so-called “Esakia spaces”.
Despite having such a good representation, the structure of free Heyting algebras is much more difficult than that of Boolean algebras or distributive lattices: already the free algebra on one generator is infinite, observed by Rieger [Rieger1952-RIEOTL] and Nishimura [nishimuralattice]. The free algebra on two generators is notoriously difficult, having no easy lattice theoretic description.
As a result, considerable attention has been devoted to describing such objects through recourse to more combinatorial and topological structures. Key results in this direction include:
@@
    The construction of the $n$-universal model (due independently to Bellissima [Bellissima1986], Grigolia [Grigolia1995]) and Shehtman [shehtmanrieger], which clarified the size and structure of irreducible elements of the free $n$-generated algebra, as well as some of its properties like completeness;
@@
    The construction of the free $n$-generated Heyting algebra by Ghilardi [ghilardifreeheyting] (which generalized previous work by Urquhart [Urquhart1973], and was in turn generalized for all finitely presented Heyting algebras by Butz [Butz1998]), which allowed to show that such algebras in fact carry more structure, like a dual connective to the implication.
Despite the importance of such advances, the situation presently in the literature presents a big contrast between what is known about Boolean algebras, distributive lattices and Heyting algebras: the former have descriptions of their free algebras, coproducts, and free one-step extensions for any number of generators, and any original algebra, whilst the latter are only available for very limited situations, involving finite algebras.
The main contribution of this paper is to provide a generalization of the constructions presented by Ghilardi to the infinite case. The main hurdle in this is realizing wherein finiteness plays a role, to allow using Priestley duality in place of Birkhoff duality for finite distributive lattices, a task which turns out to not be obvious. This turns out to amount to describing, in dual terms, the left adjoint to the inclusion of $\mathbf{HA}$, the category of Heyting algebras, in $\mathbf{DL}$, the category of distributive lattices.
Having such a description at hand, we employ such a construction for several purposes:
  1. 1.
    We provide a description of the free Heyting algebras on any number of generators;
  2. 2.
    We provide an associated concrete description of the coproducts of Heyting algebras, as well as that of the pushout of Heyting algebras. The latter is in turn used to give a new proof to the well-known fact that the category of Heyting algebras has the amalgamation property.
@@
    Using the construction of coproducts we prove a new result showing that the category of Heyting algebras is co-distributive, as well as several related properties concerning co-extensivity of this category.
We then proceed to extend these results in several directions, staying in the realm of Heyting algebras and related categories. We first show that this construction can be adapted when dealing with some subvarieties of Heyting algebras. Rather than providing a general theory, we focus on three important and illustrative cases: Boolean algebras, $\mathsf{KC}$ algebras and $\mathsf{LC}$ algebras. Finally, we look at the relationship between the category of posets with p-morphisms and the category of posets with monotone maps, and provide a right adjoint to the inclusion which is heavily inspired by the above constructions. The latter has been used in joint work with Nick Bezhanishvili [coalgebraicintuitionisticmodal] to address an open problem concerning coalgebraic representations of intuitionistic modal algebras. We also allude to further results on the category of Heyting algebras which we plan to develop in a sequel to this paper [openmapsandinterpolantcomplexity].
The structure of the paper is as follows: in Section 2 we recall some basic preliminaries, and fix some notation moving forward. Section 3 contains the main technical tools, in the form of the description of the free Heyting algebra generated by a distributive lattice preserving a given set of relative pseudocomplements; in Section 4 we showcase some basic applications of the aforementioned theory to the study of Heyting algebras: some well-known results, like the amalgamation theorem, are shown here through direct methods, as well as some seemingly new results, like the co-distributivity of the category of Heyting algebras. In Section 5 we look at the case of subvarieties of Heyting algebras and illustrate how these constructions can be fruitfully analysed there. In Section 6 we look at the corresponding relationship between the category of posets with p-morphisms and the category of posets with monotone maps. We conclude in Section 7 with some outlines of future work.
@@ 2 Preliminaries
We assume the reader is familiar with the basic theory of Boolean algebras and Distributive lattices, as well as with Stone duality (see for example [Davey2002-lr]). Throughout we will use the term “distributive lattices” for bounded distributive lattices, i.e., our lattices will always have a bottom element $(0)$ and a top element $(1)$. We also assume the reader is familiar with Heyting algebras and their elementary theory (see e.g. [Esakiach2019HeyAlg]).
Key to our results will be the use of duality theory for distributive lattices and Heyting algebras, in the form of Priestley and Esakia duality. Recall that an ordered topological space is a triple $(X,\leq,\tau)$ where $(X,\leq)$ is a poset and $(X,\tau)$ is a topological space. Often we refer to $(X,\leq)$, or just $X$ as an ordered topological space, when no confusion arises. Given such a space, and a subset $U$, we write ${\uparrow}U=\{x:\exists y\in U,x\geq y\}$ (symmetrically, ${\downarrow}U$), and say that $U$ is an upset (downset) if $U={\uparrow}U$ ($U={\downarrow}U$).
@@
Let $X$ be an ordered topological space. We say that $X$ is a Priestley space if $X$ is compact and whenever $x\nleq y$ there is some clopen upset $U$ such that $x\in U$ and $y\notin U$.
We say that a Priestley space $X$ is Esakia if whenever $U$ is clopen, ${\downarrow}U$ is clopen as well.
Definition 2.
Let $X,Y$ be Priestley spaces. We say that $f:X\to Y$ is a p-morphism if whenever $f(x)\leq y$ there is some $x^{\prime}\geq x$ such that $f(x^{\prime})=y$.
We denote by $\mathbf{DLat}$ the category of distributive lattices and lattice homomorphisms, and by $\mathbf{Pries}$ the category of Priestley spaces and monotone and continuous functions. We denote by $\mathbf{HA}$ the category of Heyting algebras and Heyting algebra homomorphisms, and by $\mathbf{Esa}$ the category of Esakia spaces and p-morphisms.
Theorem 3.
The category $\mathbf{DLat}$ is dually equivalent to the category $\mathbf{Pries}$, and the category $\mathbf{HA}$ is dually equivalent to the category $\mathbf{Esa}$.
The constructions in this duality are similar to those of Stone duality: one takes for a given distributive lattice or Heyting algebra $\mathcal{H}$ the spectrum $\mathsf{Spec}(H)=\{x:x\text{ is a prime filter in $H$}\}$; and conversely, given a Priestley or Esakia space $X$, one takes $\mathsf{ClopUp}(X)$, the set of clopen upsets. More specifically, given a distributive lattice $\mathcal{D}$, we denote the representation map as
$\varphi:\mathcal{D}\to\mathsf{ClopUp}(X_{D})$
where $\varphi(a)=\{x\in X_{D}:a\in x\}$. The reason this map is bijective follows from the prime ideal theorem: given $\mathcal{D}$ a distributive lattice, and $A\subseteq D$, we write $\mathsf{Fil}(A)$ for the filter generated by $A$, and similarly, $\mathsf{Id}(A)$ for the ideal generated by $A$. We also recall that the following prime filter theorem holds for distributive lattices:
@@
As far as morphisms work, given a distributive lattice homomorphism $f:D\to D^{\prime}$, the dual morphism is given by $f^{-1}:X_{D^{\prime}}\to X_{D}$, which maps prime filters to prime filters. Similarly, if $p:X\to Y$ is a monotone continuous map, then $p^{-1}:\mathsf{ClopUp}(Y)\to\mathsf{ClopUp}(X)$ is a distributive lattice homomorphism; this restricts to Esakia spaces and Heyting algebras in the obvious way.
A fact that we will often use below is that if $(X,\leq)$ is a Priestley space, and $Y\subseteq X$ is a closed subset, then $(Y,\leq_{\restriction Y})$ is a Priestley space as well. We also note the following important, though easy to prove, fact:
@@
Given an arbitrary subset $U\subseteq X$ of an ordered topological space, we will then write $\Box U=X-{\downarrow}X-U$.
We also recall the following construction, which is part of the folklore of the subject (see [gehrkevangoolbooktopologicalduality] for an in-depth discussion):
Definition 6.
Let $X$ be an arbitrary set. Let $\mathbf{D}(X)$ be the distributive lattice dual to the Priestley space
$F_{D}(X)\coloneqq\mathbf{2}^{X}$
where $\mathbf{2}$ is poset $0<1$, and $F_{D}(X)$ is given the product topology and the product ordering.
We then have the following well-known fact:
Proposition 7.
If $X$ is an arbitrary set, then $\mathbf{D}(X)$ is isomorphic to the free distributive lattice on $X$ many generators. Moreover, if $X$ is finite, then $F_{D}(X)$ is isomorphic to the discrete topology on on $(\mathcal{P}(X),\supseteq)$.
We will also occasionally need the Rieger-Nishimura ladder, which is the Esakia dual of the $1$-generated Heyting algebra. This can be constructed (see e.g. [Bezhanishvili2006]) in the following way: starting with three points $2,1,0$, and order them by setting $2<1$, and let $X_{0}=\{2,1,0\}$. We then construct $X_{n+1}$ by picking for each antichain of points in $X_{n}$ $x_{0},...,x_{n}$ such that there is no point $y$ immediately below each $x_{i}$; as it turns out, one can show that $X_{n+1}$ always adds only two points. We let $X(1)=\bigcup_{n\in\omega}X_{n}\cup{\infty}$, and order $\infty$ below every other point; we then give $X(1)$ the topology based on finite sets and cofinite sets containing $\infty$. The resulting space is depicted in Figure
[…]
Figure 1: Rieger-Nishimura Ladder
@@
Let $(\mathcal{D}_{n},\iota_{i,j})$ be a chain of distributive lattices, connected by embeddings $i_{i,j}:\mathcal{D}_{i}\to\mathcal{D}_{j}$ for $i<j$ satisfying the usual compatibility laws of directed systems. Let $(X_{n},p_{i,j})$ be the inverse system of Priestley spaces one obtains by dualising all arrows. Then the directed union of the directed chain, $\mathcal{D}_{\omega}$, is dual to the projective limit $X_{\omega}$ of the inverse system. Moreover, if each $X_{n}$ is an Esakia space, and each $p_{i,j}$ is a p-morphism, then $X_{\omega}$ is likewise an Esakia space.
We also recall the well-known construction of the Vietoris hyperspace, due to Leopold Vietoris [Vietoris1921] (see also [guramremarksonhyperspacesforpriestleyspaces]): let $(X,\leq)$ be a Priestley space. Let $V(X)$ be the set of all closed subsets of $X$. We give this set a topology by considering a subbasis consisting, for $U,V\subseteq X$ clopen sets:
@@
This is sometimes called the “hit-and-miss topology”, and the resulting space is called the Vietoris hyperspace. On this space we define an order relation $C\preceq D$ if and ony if [^1] The reason this order is reverse inclusion has to with the choice to use upsets rather than downsets; in [ghilardifreeheyting], the author uses downsets, which is why the swap no reversion is needed. $D\subseteq C$. Then we have the following:
@@
The fact that $V(X)$ is compact is a standard fact (see e.g. [Engelking1989-qo]), but we prove it here for completeness. assume that $V(X)=\bigcup_{i\in I}[U_{i}]\cup\bigcup_{j\in J}\langle V_{j}\rangle$ is a cover by clopen sets. Now consider $C=X-\bigcup_{i\in I}V_{i}$. If $C$ is empty, then the sets $V_{i}$ cover $X$ and so finitely many of them cover $X$, say $V_{i_{0}},...,V_{i_{n}}$. Then if $A\in V(X)$, then $A$ is non-empty and so contains some $x$ which must lie in some $V_{i_{k}}$ for $k\in\{1,...,n\}$, and so $A\in\langle V_{i_{k}}\rangle$. So the sets $\{\langle V_{i_{k}}\rangle:k\in\{1,...,n\}\}$ form a finite subcover.